OpenStudy (anonymous):

Calc help . Find the absolute extrema of the fn over the closed interval. f(x)=x/(sqr.rt of x^2+1) [0, 2]

OpenStudy (anonymous):

\[f(x)=\frac{x}{\sqrt{x^2+1}}\]?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

did you take the derivative?

OpenStudy (anonymous):

yes thats what i did

OpenStudy (anonymous):

did you get \[f'(x)=\frac{1}{\sqrt{x^2+1}^3}\]?

OpenStudy (anonymous):

nope :[

OpenStudy (anonymous):

oh too bad

OpenStudy (anonymous):

perhaps you got a big fat mess that looked hard to simplify

OpenStudy (anonymous):

in the end i got x^2+1/(x^2+1)^1/2

OpenStudy (anonymous):

you need the quotient rule for this one

OpenStudy (anonymous):

i think so :( hey that was mean :`(

OpenStudy (anonymous):

that wat i did

OpenStudy (anonymous):

ok lets do that second, first lets answer the question the derivative is always positive, which means the function is always increasing therefore the min is at the left endpoint, namely at \(x=0\) and the max is at \(x=2\)

OpenStudy (anonymous):

ok :l

OpenStudy (anonymous):

for the derivative you need \[\frac{gf'-fg'}{g^2}\] with \[f(x)=x, f'(x)=1,g(x)=\sqrt{x^2+1}, g'(x)=\frac{x}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you get by direct substitution \[\frac{\sqrt{x^2+1}-\frac{x^2}{\sqrt{x^2+1}}}{x^2+1}\]

OpenStudy (anonymous):

huh? whatha :`(

OpenStudy (anonymous):

now to simplify this ugly mess, multiply top and bottom by \(\sqrt{x^2+1}\) to clear the compound fraction

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

uh oh maybe we should go back and make sure it is clear how to find the derivative using the quotient rule, unless what i wrote was clear i sense it was not

OpenStudy (anonymous):

wait ok dont u h=get -2x^2

OpenStudy (anonymous):

no i get it its just that i got -2x^2 instead of -x^2 :l

OpenStudy (anonymous):

no if you multiply top and bottom by \(\sqrt{x^2+1}\) the numerator becomes \[x^2+1-x^2\]

OpenStudy (anonymous):

i mean for the numerator i got okay i got (x^2+1)^1/2-x^2/2(x^2+1)^1/2 all over x^2+1

OpenStudy (anonymous):

\[\left(\frac{\sqrt{x^2+1}-\frac{x^2}{\sqrt{x^2+1}}}{x^2+1}\right)\times \frac{\sqrt{x^2+1}}{\sqrt{x^2+1}}\] \[=\frac{x^2+1-x^2}{(x^2+1)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

yes what you wrote is correct

OpenStudy (anonymous):

okay :}

OpenStudy (anonymous):

this is right (x^2+1)^1/2-x^2/2(x^2+1)^1/2 all over x^2+1

OpenStudy (anonymous):

so then.....where du i go from what i had multiply by (x^2+1)^1/2??

OpenStudy (anonymous):

but in the end, after getting rid of the compound fraction you are left with \[f'(x)=\frac{1}{(x^2+1)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

yes that is the easiest way to do it

OpenStudy (anonymous):

o okay :]

OpenStudy (anonymous):

the radical is gone from the numerator, leaving only 1 up top since the numerator is 1, it is never zero, so there are no critical points the derivative is always positive function is always increasing here is a nice picture http://www.wolframalpha.com/input/?i= \frac{x}{\sqrt{x^2%2B1}}+domain+0..2

OpenStudy (anonymous):

link didn't work but if you copy and paste you will see it

OpenStudy (anonymous):

ok i sorta got it thanks!

OpenStudy (anonymous):

like i said (not trying to be mean, really) it is the messy algebra that is a pain only reason i know it is because i have seen it several times